1. What mass of natural gas (CH4) must you burn to emit 260 kJ of heat?

CH4(g)+2O2(g)→CO2(g)+2H2O(g)

ΔH∘rxn=−802.3kJ

2.Consider the following thermochemical equation for the combustion of acetone, C3H6OC3H6O, the main ingredient in nail polish remover:

C3H6O(l)+4O2(g)→3CO2(g)+3H2O(g)ΔH∘rxn=−1790kJ

If a bottle of nail polish remover contains 173 mL of acetone, how much heat would be released by its complete combustion? The density of acetone is 0.788 g/mL

3.Zinc metal reacts with hydrochloric acid according to the following balanced equation:

Zn(s)+2HCl(aq)→ZnCl2(aq)+H2(g)

When 0.119 gg of Zn(s)Zn(s) is combined with enough HCl to make 51.6 mL of solution in a coffee-cup calorimeter, all of the zinc reacts, raising the temperature of the solution from 21.6 ∘C to 24.6 ∘C

Find ΔHrxn for this reaction as written. (Use 1.00 g/mL for the density of the solution and 4.18 J/(g⋅∘C) as the specific heat capacity.)

Express your answer in kilojoules per mole to two significant figures.
Answer #1

1) Given reaction:

CH4 + 2O2 --------------------> CO2 + 2 H2O ; ∆H°rxn = -802.3 KJ

Since, 1 mole of CH4 emits 802.3 KJ energy

So, moles of CH4 required to emit 260 KJ energy

= 260 / 802.3 = 0.324 moles

So mass of natural gas (CH4) required

= moles× molar mass

= 0.324 × 16 = 5.185 grams (Answer)